alex is x years old. June is 7 years older than alex. In 5 years, their total combined age will be 31 years
Write a linear equation for their total combined age in 5 years

Answer by ankor@dixie-net.com(22740)

Write an equation for each statement
alex is x years old.
a = x
June is 7 years older than alex.
j = x+7
In 5 years, their total combined age will be 31 years
(x+5) + (j+5) = 31
Write a linear equation for their total combined age in 5 years
Replace j with (x+7)
x + 5 + (x+7) + 5 = 31
2x + 17 = 31
we can solve it
2x = 31 - 17
2x = 14
x = 7 is A's age
then
14 = J's age